# Integrated Sensing and Communications Over the Years: An Evolution Perspective

### v2

@@ I-A Background and Motivation
Wireless communication networks are widely regarded as foundational and strategic industries, serving as the backbone for emerging sectors such as the industrial Internet of Things (IIoT) and intelligent manufacturing [1]. Their continued evolution is considered a critical element in global science and technology agendas. Since the advent of radio systems in 1934, the independent development of communication and radar systems has introduced several challenges, including inefficient and competitive use of wireless resources [2], duplication of infrastructure and hardware [3], and limited cross-system interoperability [4]. Given that electromagnetic signals inherently support both environmental sensing and data transmission, maintaining separate system designs has become increasingly redundant.
@@
to Multi-Modal Security and Privacy Standardization Processes
[5] $\times$ $\times$ $\times$ $\times$ IoT applications
[6] $\times$ $\times$ $\times$ Dual-functional wireless networks in 6G
[7] $\times$ $\times$ $\times$ $\times$ Perceptive mobile network structure, signal processing
[8] $\times$ $\times$ $\times$ $\times$ Resource, node, and infrastructure-level cooperation
[9] $\times$ $\times$ $\times$ Vehicular networks, 6G applications
@@
[11] $\times$ $\times$ $\times$ Multi-view sensing architectures and designs
[12] $\times$ $\times$ $\times$ $\times$ ISAC hardware testbeds and performance evaluation
[13] $\times$ $\times$ $\times$ $\times$ Signal Processing
[14] $\times$ $\times$ $\times$ $\times$ Wi-Fi sensing using CSI
[15] $\times$ $\times$ $\times$ Historical development of ISAC
@@
[17] $\times$ $\times$ $\times$ $\times$ Spectrum and hardware sharing
[18] $\times$ $\times$ $\times$ $\times$ Waveform design and MIMO processing
[19] $\times$ $\times$ $\times$ $\times$ UAV network structure
[20] $\times$ $\times$ $\times$ Multi-carrier waveform design and learning techniques
@@
[22] $\times$ AI/ML-assisted ISAC architecture and secure implementation
[23] $\times$ $\times$ Visionary techniques for 6G ISAC
@@
[25] $\times$ Secure architecture across HW/Omniscient/Application layers
@@
[27] $\times$ $\times$ Unified multi-modal ISAC framework
@@
[30] $\times$ $\times$ $\times$ RIS for multi-function ISAC
@@
In response, integrated sensing and communication (ISAC) has emerged since the mid-2010s as a unified paradigm in which wireless systems are designed to simultaneously sense the physical environment and transmit information through the co-design of shared waveforms, spectrum, and hardware platforms [5, 6]. In this context, sensing refers broadly to radar-type functions, including object detection, tracking, and recognition, as well as behavior analysis and environmental awareness. Since its formal introduction, and particularly after 2020, significant progress has been achieved in key areas such as joint signal design, multifunctional antenna systems, spectrum sharing, and edge intelligence. Indeed, ISAC is transitioning from research to large-scale deployment, driven by both industrial adoption and standardization initiatives. In the industrial domain, Huawei has advanced ISAC deployment in vehicular networks [31], Ericsson has extended its application to intelligent transportation and urban infrastructure [32], and Qualcomm has explored millimeter-wave ISAC for autonomous driving and IIoT applications [33]. On the standardization side, 3GPP has identified 32 ISAC use cases for 6G [34], ETSI’s Industry Specification Group has proposed 18 additional cases [35], and IEEE finalized the first Wi-Fi sensing standard (802.11bf) in October 2024. At the global policy level, the ITU has designated ISAC as one of six core 6G application scenarios [36], and the World Economic Forum has recognized it among the top ten emerging technologies of 2024 [37]. Collectively, these developments underscore ISAC’s accelerating momentum and its potential to become a cornerstone technology in next-generation (6G) wireless networks.
@@ I-B Related Survey Papers
Despite substantial progress, a comprehensive survey capturing the full breadth of ISAC research remains lacking, as existing works primarily address specific technical dimensions or application domains. Prior surveys have examined spectrum and hardware sharing [17], waveform and multiple-input multiple-output (MIMO) techniques [18], and Internet of Things (IoT)-centric implementations [5]. Signal processing approaches for 5G-Advanced and 6G have also been discussed [13, 38], while channel state information (CSI)-based Wi-Fi sensing [14], along with historical and theoretical foundations [15, 16], has established the conceptual basis for subsequent advances. Building on these foundations, research has progressively expanded toward enabling technologies across the physical, deployment, and system levels. At the waveform and modeling level, efforts focus on multi-carrier and dual-functional signaling, ISAC channel modeling, and interference management in dense networks [20, 6, 26, 28]. At the deployment layer, reconfigurable intelligent surfaces (RIS/IRS) have been extensively studied for cooperative networking [21, 39, 40], with extensions to RIS-assisted non-orthogonal multiple access (NOMA) to enhance spectrum and sensing efficiency [24]. Application-driven studies have addressed vehicular network deployments [9] and UAV-based implementations [19], while the notion of perceptual mobile networks has been introduced [10] to envision infrastructures capable of dynamic mobile sensing and communication (S&C). Complementing theoretical progress, hardware testbeds and empirical evaluations have been developed to validate ISAC concepts in practice [12]. Beyond these, integration with artificial intelligence (AI) and security has enabled adaptive transmission, environmental inference, and secure IoT architectures [22, 29, 25]. More recently, the trajectory toward multi-modal perception has been marked by synesthesia-inspired frameworks, multi-view designs, and generalized system principles [27, 11, 23, 7].
Viewed holistically, these studies outline a broad but fragmented landscape, demonstrating steady progress toward increasingly integrated, intelligent, and multifaceted ISAC systems. As summarized in Table I-A, however, a unified perspective that connects spectrum evolution, network scaling, sensing modalities, system-level challenges, and standardization remains absent. This work addresses the gap by presenting a multi-dimensional evolutionary analysis of ISAC, examining the expansion of operating spectrum from RF to optical bands, the scaling of network architectures from single-cell to multi-cell, and the transition from single-modal to multi-modal sensing empowered by edge intelligence. It further considers emerging security and privacy requirements alongside ongoing standardization activities, thereby providing a coherent view of the state of the art and highlighting open challenges for 6G implementation.
@@ III-A2 Unified Design
where $f_{0}$ is the central frequency and $k$ denotes the chirp rate. The linear frequency sweep provides fine range resolution and robustness to noise. Compared with OFDM, FMCW generally requires much lower sampling rates, which simplifies receiver processing. However, this advantage comes at the expense of significantly reduced data transmission rates, making FMCW more suitable for sensing-centric designs [115].
Additionally, index modulation schemes allow information to be conveyed through antenna or frequency selection without altering the primary radar waveform. Recent studies have explored these embedding and modulation techniques, emphasizing trade-offs in symbol rate and beam fidelity. For example, [116] demonstrated a spatial modulation-based ISAC prototype achieving a symbol error rate (SER) below $10^{-2}$ at 15 dB SNR while maintaining angular accuracy within $1.8^{\circ}$. Similarly, [99] proposed a joint beamforming framework for multi-user MIMO ISAC, maintaining radar gain degradation within 3 dB while achieving near-optimal communication sum rate. However, sensing-centric designs generally offer limited communication capacity and impose higher receiver processing complexity.
@@
Conversely, communication-centric designs superimpose sensing functions onto existing communication waveforms, most notably OFDM and orthogonal time–frequency space (OTFS) [117] modulation [118, 119].
@@
In contrast to separate optimization strategies, joint S&C design seeks to balance trade-offs through mathematical formulation. By incorporating performance metrics from both domains as joint objectives, this approach addresses inherent conflicts using advanced convex formulations and emphasizes unified performance criteria.
Several metrics have been widely adopted in joint waveform design. Among them, mutual information (MI) serves as a versatile criterion for both sensing and communication. It captures information efficiency in data transmission and target detection, and remains effective under imperfect channel conditions such as partial CSI or clutter.
@@
where $\rho\in[0,1]$ is a weighting factor, I denote the MI, and $Y_{c}$, $Y_{s}$, $H_{c}$, and $H_{s}$ represent the received signals and channel/target impulse responses for communication and sensing, respectively. This framework allows for identifying the Pareto optimal MI frontier, effectively illustrating the trade-offs between S&C in ISAC systems.
Given that resources such as power and bandwidth are shared in ISAC systems, a unified resource allocation framework is required. The objective is to prioritize critical sensing tasks, such as detecting high-risk targets, while still guaranteeing communication requirements. This consideration motivates the following optimization problem
@@
where the achievable sum-rate $R_{\mathbf{p},\mathbf{b}}$ is required to exceed a threshold $R$, the power allocation vector $\mathbf{p}$ and bandwidth allocation vector $\mathbf{b}$ are assigned across $M$ objects under total resource constraints $P_{\text{total}}$ and $B_{\text{total}}$, respectively. Typical sensing QoS functions are defined as follows [129]:
@@
    where $P_{D}$ denotes the probability of target detection, and $\epsilon_{q}$ is the normalized sensing channel gain associated with transmit power $p_{q}$ and array gain.
@@
    where the CRB provides the lower bound on the estimation errors of the distance $d_{q}$ and the angle $\theta_{q}$. These bounds are inversely proportional to the transmit power $p_{q}$ and, in the case of distance estimation, also to the allocated bandwidth $b_{q}$. Since $b_{q}$ is simultaneously allocated for communications, interference mitigation, and optimal resource utilization are essential.
@@
    where the PCRB is derived by recursively computing the Fisher information matrix for the target state, thereby reflecting dynamic tracking accuracy. The transmit power $p_{q,n}$ and bandwidth $b_{q,n}$ are dynamically allocated in each time slot $n$ to satisfy real-time communication requirements, such as latency constraints.
@@
Recent studies have introduced frameworks to address joint optimization challenges. The closed-form performance characterization of ISAC metrics such as probability of false alarm and probability of detection under power allocation was derived in [130], offering direct insights into how transmit power impacted sensing accuracy. A flexible allocation strategy was introduced in [131], enabling efficient distribution of limited power and bandwidth between S&C. Furthermore, an optimal solution based on the sum-of-ratios method was presented in [132], coordinating transmit power, time allocation, and quantization levels across distributed devices to mitigate computational distortion.
These contributions collectively highlight the importance of efficient, adaptive, and task-specific joint resource allocation, laying the theoretical groundwork for large-scale ISAC deployment.

### v3

@@ I-A Background and Motivation
Wireless communication networks are widely regarded as foundational and strategic industries, serving as the backbone for emerging sectors such as the industrial Internet of Things (IIoT) and intelligent manufacturing [cui2025integratedsensingcommunicationmultifunctional]. Their continued evolution is considered a critical element in global science and technology agendas. Since the advent of radio systems in 1934, the independent development of communication and radar systems has introduced several challenges, including inefficient and competitive use of wireless resources [10255711], duplication of infrastructure and hardware [zheng2019radar], and limited cross-system interoperability [8828023]. Given that electromagnetic signals inherently support both environmental sensing and data transmission, maintaining separate system designs has become increasingly redundant.
@@
to Multi-Modal Security and Privacy Standardization Process
[9606831] $\times$ $\triangle$ $\times$ $\times$ $\times$ IoT applications
[9737357] $\times$ $\triangle$ $\times$ $\times$ $\triangle$ Dual-functional wireless networks in 6G
[9585321] $\times$ $\triangle$ $\times$ $\times$ Perceptive mobile network structure, signal processing
[wei2024deep] $\times$ $\times$ $\times$ $\times$ Resource, node, and infrastructure-level cooperation
[10418473] $\times$ $\triangle$ $\times$ $\times$ $\triangle$ Vehicular networks, 6G applications
@@
[10049304] $\times$ $\triangle$ $\times$ $\times$ Multi-view sensing architectures and designs
[10049817] $\times$ $\triangle$ $\times$ $\times$ $\times$ ISAC hardware testbeds and performance evaluation
[8828030] $\times$ $\triangle$ $\times$ $\times$ $\times$ Signal processing for ISAC
[10.1145/3310194] $\times$ $\triangle$ $\times$ $\times$ $\times$ Wi-Fi / CSI-based sensing
[10188491] $\times$ $\triangle$ $\times$ $\times$ $\triangle$ Historical development of ISAC
@@
[8999605] $\triangle$ $\times$ $\times$ $\times$ $\times$ Spectrum and hardware sharing
[9540344] $\times$ $\triangle$ $\times$ $\times$ $\times$ Waveform design and MIMO processing
[10004900] $\times$ $\triangle$ $\times$ $\times$ $\times$ UAV network structure
[10770016] $\times$ $\triangle$ $\triangle$ $\times$ $\times$ Multi-carrier waveform design and learning techniques
@@
[11015430] $\times$ $\triangle$ $\triangle$ $\triangle$ AI/ML-assisted ISAC architecture and secure implementation
[10536135] $\triangle$ $\triangle$ $\times$ $\times$ $\triangle$ Visionary techniques for 6G ISAC
@@
[10608156] $\triangle$ $\triangle$ $\times$ Secure architecture across HW/Omniscient/Application layers
@@
[10330577] $\triangle$ $\times$ $\times$ Unified multi-modal ISAC framework
@@
[10833623] $\triangle$ $\times$ $\times$ $\times$ RIS for multi-function ISAC
@@
(*) ✓   Indicates major and systematic coverage of the corresponding dimension, typically supported by dedicated analysis or structured discussion. (*) $\triangle$   Indicates partial or incidental coverage, where the dimension is mentioned but not treated as a primary focus. (*) $\times$   Indicates that the dimension is not explicitly addressed in the survey.
In response, integrated sensing and communication (ISAC) has emerged since the mid-2010s as a unified paradigm in which wireless systems are designed to simultaneously sense the physical environment and transmit information through the co-design of shared waveforms, spectrum, and hardware platforms [9606831, 9737357]. In this context, sensing refers broadly to radar-type functions, including object detection, tracking, and recognition, as well as behavior analysis and environmental awareness. Since its formal introduction, and particularly after 2020, significant progress has been achieved in key areas such as joint signal design, multifunctional antenna systems, spectrum sharing, and edge intelligence. Indeed, ISAC is transitioning from research to large-scale deployment, driven by both industrial adoption and standardization initiatives. In the industrial domain, Huawei has advanced ISAC deployment in vehicular networks [huawei2022isac], Ericsson has extended its application to intelligent transportation and urban infrastructure [ericsson2022isac], and Qualcomm has explored millimeter-wave ISAC for autonomous driving and IIoT applications [qualcomm2022isac]. On the standardization side, 3GPP has identified 32 ISAC use cases for 6G [3gpp2022isac], ETSI’s Industry Specification Group has proposed 18 additional cases [ESTI2025isac], and IEEE finalized the first Wi-Fi sensing standard (802.11bf) in October 2024. At the global policy level, the ITU has designated ISAC as one of six core 6G application scenarios [itu2021vision], and the World Economic Forum has recognized it among the top ten emerging technologies of 2024 [wef2024emergingtech]. Collectively, these developments underscore ISAC’s accelerating momentum and its potential to become a cornerstone technology in next-generation (6G) wireless networks.
@@ I-B Related Survey Papers
A broad collection of surveys has examined ISAC from complementary perspectives. Early contributions focused on spectrum and hardware reuse mechanisms [8999605], as well as waveform and multiple-input multiple-output (MIMO) techniques [9540344]. These works were followed by surveys on internet of things (IoT)-oriented ISAC architectures [9606831], signal-processing techniques for 5G-Advanced/6G [8828030, 10012421], channel state information (CSI)-based Wi-Fi and RF sensing [10.1145/3310194], and the historical and theoretical foundations of joint radar–communication systems [10188491, 9705498], forming the foundations of ISAC research.
Built on these foundations, subsequent surveys have expanded ISAC along several technically distinct directions. One important stream has delved into physical-layer mechanisms, including multi-carrier and dual-functional waveform construction, ISAC-tailored channel modeling, and interference-management approaches for dense or heterogeneous deployments [10770016, 9737357, 10646523, 10770127]. These studies deepen the understanding of signal construction, propagation behavior, and hardware-related tradeoffs in practical ISAC deployments. At the deployment and architectural level, reconfigurable intelligent surfaces (RIS/IRS) have been widely investigated for cooperative networking [10726912, 10702570, 10833779], with further extensions toward RIS-assisted non-orthogonal multiple access (NOMA) for enhanced joint sensing and communication (S&C) efficiency [10781415]. Application-driven surveys have addressed ISAC in vehicular networks [10418473] and uncrewed aerial vehicles (UAV)-based implementations [10004900], while the notion of perceptual mobile networks has been introduced to describe mobile infrastructures capable of dynamic S&C fusion [9296833]. Complementing analytical advances, hardware testbeds and empirical evaluations have validated ISAC concepts in practice [10049817].
Beyond physical-layer and deployment-oriented studies, another stream of literature explores ISAC’s integration with artificial intelligence (AI), security, and emerging perception paradigms. These efforts span adaptive and environment-aware transmission, secure IoT architectures, and data-centric learning frameworks for ISAC [11015430, 10989512, 10608156]. Recent studies on multi-view sensing, generalized architectural principles, and emerging methods for multi-modal S&C integration [10330577, 10049304, 10536135, 9585321] highlight the ongoing shift toward richer environmental awareness and learning-driven ISAC capabilities.
Viewed collectively, these surveys offer substantial depth across specific ISAC dimensions—waveforms, channel modeling, network cooperation, RIS-enhanced operation, multi-modal perception, security, and AI-assisted inference. However, their viewpoints remain inherently theme-centered: each provides a focused treatment on a subset of these dimensions, typically around a single dominant theme (e.g., security-centric, RIS-centric, or multi-modal-centric), but none attempts to synthesize how ISAC has evolved simultaneously across spectrum expansion, network-scale development, sensing-modality diversification, system-level optimization and security constraints, and progressing standardization.
As summarized in Table I-A, the existing surveys form a collection of rich but isolated slices of the broader ISAC landscape. A consolidated multi-dimensional evolutionary perspective that connects spectrum expansion, network-scale development, sensing-modality diversification, system-level optimization and standardization into a coherent trajectory is still not fully developed, even as such a perspective becomes increasingly important for guiding ISAC toward heterogeneous, scalable, and deployment-ready 6G systems.
@@ I-C Contributions
Taken together, these contributions deliver a unified evolutionary framework that connects previously disjointed research themes into a coherent system-level narrative. This multi-dimensional synthesis, spanning spectrum, architecture, perception, security, and standardization, fills a key gap in existing ISAC surveys and lays the foundation for the broader adoption of ISAC within heterogeneous, scalable, and deployment-ready 6G systems.
@@ II Operating Spectrum across RF and Optical
II-A RF-based ISAC
RF technology, operating in the 3 kHz–300 GHz range, remains the foundation of wireless communication and sensing. In ISAC systems, transmitter design begins with stable carrier generation via oscillators and synthesizers, followed by modulation through mixers and upconverters to embed data or probing signals. The RF chain, comprising low-noise amplifiers (LNAs), power amplifiers (PAs), and RF filters, conditions the signal before impedance matching and directional coupling delivers it to the antenna for electromagnetic radiation. In receiver design, incident signals are captured by the antenna, amplified, downconverted, and digitized through analog-to-digital converters (ADCs) [walden2002analog], followed by digital signal processing (DSP) to recover transmitted data or extract environmental features [sturm2011waveform].
While antenna arrays, particularly massive MIMO and millimeter-wave (mmWave) configurations, enable beamforming, beam steering, and spatial multiplexing for concurrent multi-user communication and sensing, overall performance is equally constrained by RF front-end and baseband hardware. Current ISAC research emphasizes antenna architectures, suggesting a future need for deeper hardware-level performance analysis that holistically optimizes LNAs, PAs, RF filters, ADCs, and DSP modules to maximize joint S&C efficiency.
II-A1 Advanced Antenna Technologies
Leveraging spatial degrees of freedom to enhance both communication and sensing, multi-antenna technologies constitute a key enabler for next-generation ISAC. Three architectural paradigms are especially impactful: centralized arrays, distributed arrays, and movable/fluid arrays.
Centralized Antenna Architectures
These encompass compact and sparse array deployments. Compact arrays, with closely packed elements, offer high beamforming gain and spatial multiplexing capability, but at the cost of increased hardware complexity and a limited field of view. In contrast, sparse arrays employ non-uniform geometries (e.g., nested, coprime [5609222]) to synthesize large virtual apertures, reducing the number of active RF chains while retaining beamforming functionality. Although resolution and robustness depend on array design and signal processing complexity, sparse architectures strike an effective balance between cost and angular resolution, making them attractive for ISAC hardware efficiency. For instance, Zeng et al. [zeng2024fixedmovableantennatechnology] derived unified Cramér–Rao bound (CRB) and signal-to-noise ratio (SNR) scaling laws for fixed and mechanically movable uniform linear arrays (ULAs), showing that doubling the aperture halves the bearing-error variance but proportionally narrows the allowable carrier-phase mismatch. Li et al. [li2025sparse] demonstrated that uniform sparse and coprime layouts, when constrained to the same aperture as dense ULAs, preserved main-lobe width and even outperformed them in sum-rate and direction of arrival (DOA) estimation root mean square error (RMSE) at sub-0 dB SNR, using only one-quarter of the RF chains. Building on these insights, Elbir et al. [10371037] introduced deep learning-assisted hybrid beamforming that reduced the antenna selection search space by nine orders of magnitude while limiting spectral efficiency loss to under 5% across 0–20 dB SNR. Taken together, these studies underscore that sparse-geometry optimization, combined with learning-based beamforming, can achieve sub-degree angular estimation accuracy with substantially reduced hardware overhead.
@@ III-A2 Unified Design
where $f_{0}$ is the central frequency and $k$ denotes the chirp rate. The linear frequency sweep provides fine range resolution and robustness to noise. Compared with OFDM, FMCW generally requires much lower sampling rates, which simplifies receiver processing. However, this advantage comes at the expense of significantly reduced data transmission rates, making FMCW more suitable for sensing-centric designs [sturm2011waveform].
Additionally, index modulation schemes allow information to be conveyed through antenna or frequency selection without altering the primary radar waveform. Recent studies have explored these embedding and modulation techniques, emphasizing trade-offs in symbol rate and beam fidelity. For example, [9345999] demonstrated a spatial modulation-based ISAC prototype achieving a symbol error rate (SER) below $10^{-2}$ at 15 dB SNR while maintaining angular accuracy within $1.8^{\circ}$. Similarly, [liu2020joint] proposed a joint beamforming framework for multi-user MIMO ISAC, maintaining radar gain degradation within 3 dB while achieving near-optimal communication sum rate. However, sensing-centric designs generally offer limited communication capacity and impose higher receiver processing complexity.
@@
Conversely, communication-centric designs superimpose sensing functions onto existing communication waveforms, most notably OFDM and orthogonal time–frequency space (OTFS) [8288677] modulation [wei2021orthogonal, 10638525].
@@
In contrast to separate optimization strategies, joint S&C design formulates the objectives of the two functions within a unified mathematical framework. This formulation enables a direct characterization of their coupled behaviors and provides a principled view of how sensing and communication interact over shared wireless resources.
The motivation for unified formulations arises from the fundamental mismatch between the waveform requirements of the two functionalities. Sensing typically benefits from deterministic and energy-concentrated signals that enhance estimation precision[he2012waveform], whereas communication favors random and spectrally efficient signaling to maximize data throughput and robustness[cover1999elements]. If optimized independently, these different requirements often yield incompatible designs, thereby motivating joint objectives that explicitly encode the shared resource constraints and physical coupling[9737357, 10418473, 9585492].
Among various unified metrics, MI is widely adopted because it simultaneously characterizes the recoverable information in communication and the observability of target-induced perturbations required for sensing. MI is also grounded in information-theoretic limits, making it resilient to practical impairments such as partial channel knowledge, interference, and clutter—conditions frequently encountered in ISAC deployments[10466681].
@@
where $X$ denotes the jointly designed transmit signal, $\rho\in[0,1]$ is a weighting factor, I denote the MI, and $Y_{c}$, $Y_{s}$, $H_{c}$, and $H_{s}$ represent the received signals and channel/target impulse responses for communication and sensing, respectively. This formulation identifies the Pareto-optimal MI frontier, illustrating the fundamental trade-offs between S&C.
Given that power and bandwidth are consumed by sensing and communication, ISAC systems inherently require a unified resource allocation framework. Sensing accuracy improves with transmit energy and bandwidth, whereas communication simultaneously requires rate, latency, and reliability guarantees. Because these demands originate from the physical coupling of a dual-functional transceiver, resource allocation emerges as a central design challenge.
@@
where $R_{\mathbf{p},\mathbf{b}}$ denotes the achievable communication sum-rate, required to exceed a minimum threshold $R$. The vectors $\mathbf{p}$ and $\mathbf{b}$ represent the power and bandwidth allocated across $M$ sensing targets or communication users, subject to total resource constraints $P_{\text{total}}$ and $B_{\text{total}}$, respectively.
The sensing objectives originate from classical radar detection and estimation theory [van2002optimum], forming the fundamental limits for detecting, localizing, and tracking objects under noisy reflections. Hence, these criteria appear consistently in ISAC optimization.
@@
    where $P_{D}$ denotes the probability of correctly declaring the presence of a target; $\epsilon_{q}$ is the normalized sensing channel gain associated with transmit power $p_{q}$ and array gain; $\delta$ is the detection threshold; $N_{r}$ is the number of receive antennas; and $\sigma_{r}^{2}$ is the noise variance. Physically, $P_{D}$ measures the system’s ability to detect a target under noise and clutter. A higher $P_{D}$ directly reduces miss detection in safety-critical applications (e.g., collision avoidance or intrusion monitoring), which explains why power allocation in low-SNR regimes often prioritizes reliable detection.
@@
    The CRB specifies the minimum achievable variance of unbiased estimators for range $d_{q}$ and angle $\theta_{q}$. These bounds decrease with transmit power $p_{q}$, in the case of range estimation, with allocated bandwidth $b_{q}$. The constants $\beta_{1}$ and $\beta_{2}$ depend on waveform parameters and array geometry. A larger bandwidth $b_{q}$ improves delay (range) resolution, whereas a higher transmit power $p_{q}$ enhances angular resolution through improved SNR. The composite metric $\rho_{q}$ thus reflects the achievable localization accuracy of an ISAC transceiver and is essential for beam alignment, user positioning, and environment reconstruction.
@@
    where $\xi_{q,n}$ is the dynamic target state (e.g., position and velocity), and $\mathbf{J}_{\xi_{q,n}}$ is the corresponding Fisher information matrix. The posterior CRB (PCRB) characterizes the tightest uncertainty bound for estimating time-varying trajectories given the sensing update rate. A smaller PCRB indicates more stable and accurate tracking, which is particularly critical in high-mobility environments. Since ISAC systems must simultaneously accommodate per-slot communication constraints (e.g., latency and throughput), the transmit power $p_{q,n}$ and bandwidth $b_{q,n}$ available for sensing directly determine the achievable tracking accuracy and update rate.
Recent studies have introduced frameworks addressing joint optimization challenges. Closed-form characterizations of ISAC metrics such as false-alarm and detection probabilities under power allocation were provided in [10124135], offering insights into how transmit energy affects sensing accuracy. A flexible resource allocation strategy for jointly distributing power and bandwidth was developed in [10387517]. Moreover, an optimal sum-of-ratios-based solution was proposed in [10217150] to coordinate power, time, and quantization budgets across distributed devices. These developments demonstrate that ISAC optimization objectives arise naturally from physical constraints, estimation limits, and shared-resource coupling, rather than being imposed by arbitrary design preferences.
Practical deployment considerations further underscore the relevance of these formulations. Hardware impairments, synchronization offsets, finite-rate control signaling, and dynamic traffic loads introduce deviations from idealized assumptions, affecting both sensing and communication. Because these non-idealities influence both functionalities simultaneously, joint S&C optimization aligns closely with the practical behavior of integrated transceivers[10418473, 10273396].
Overall, the existing studies indicate the increasing importance of developing efficient and adaptive joint resource allocation mechanisms in ISAC, which provide a foundation for scaling ISAC capabilities toward larger and more heterogeneous deployments.